Admitting New Partners Who Buy an Interest and Contribute Assets

The capital accounts of Trent Henry and Tim Chou have balances of $178,500 and $128,400, respectively. LeAnne Gilbert and Becky Clarke are to be admitted to the partnership. Gilbert buys one-fifth of Henry’s interest for $41,100 and one-fourth of Chou’s interest for $28,200. Clarke contributes $43,600 cash to the partnership, for which she is to receive an ownership equity of $43,600.

a1. Journalize the entry to record the admission of Gilbert. For a compound transaction, if an amount box does not require an entry, leave it blank.

a2. Journalize the entry to record the admission of Clarke.

b. What are the capital balances of each partner after the admission of the new partners?

Partner Capital Balance
Trent Henry $
Tim Chou $
LeAnne Gilbert $
Becky Clarke $

Admitting New Partner With Bonus

Cody Jenkins and Lacey Tanner formed a partnership to provide landscaping services. Jenkins and Tanner shared profits and losses equally. After all the tangible assets have been adjusted to current market prices, the capital accounts of Cody Jenkins and Lacey Tanner have balances of $37,000 and $48,000, respectively. Valeria Solano has expertise with using the computer to prepare landscape designs, cost estimates, and renderings. Jenkins and Tanner deem these skills useful; thus, Solano is admitted to the partnership at a 30% interest for a purchase price of $23,000.

a. Determine the recipient and amount of the partner bonus.
$  

b. Provide the journal entry to admit Solano into the partnership. For a compound transaction, if an amount box does not require an entry, leave it blank.

c. Why would a bonus be paid in this situation?

Apparently, Jenkins and Tanner value   offered by Solano.
Answer #1

1)

a1
Trent Henry, capital 35700 =178500/5
Tim Chou, Capital   32100 =128400/4
      LeAnne Gilbert, Capital 67800
a2
Cash 43600
     Becky Clarke, Capital   43600
b
Partner Capital Balance
Trent Henry 142800

=178500-35700

Tim Chou 96300 =1284 00-32100
LeAnne Gilbert 67800
Becky Clarke 43600

2)

Ratio of profit between jenkin and tanner = 1:1

Total capital after new capital introduced by Solano = $37,000 + $48,000 + $23,000 = $108,000

Solano share in Partnership = 30%

Therefore required share of capital by Solano = 108000 * 30% = $32,400

As capital introduced by Solano is lesser by $9,400 therefore old partner capital share will be given to Solano in their profit sharing ratio i.e. 1:1

Therefore recepient of bonus is solano and amount of partner bonus = $9,400

Solution b:

Journal Entries
Event Particulars Debit Credit
a Cash Dr $23,000.00
Jenkins Capital Dr $4,700.00
Tanner's Capital Dr $4,700.00
      To Solano's Capital $32,400.00
(To record solano admission in partnership)

Solution c:

The bonus is paid in this transaction becaase apparently, jenkins and tanner value is higher than value offered by Solano.